Forces in a Three-Charge System Coulombs law for the magnitude of the force F between two particles with charges Q and Q separated by a distance d is where K 1 and ATeo eo 8.854 x 10 12 C/(N ma) s the permittivity of free space. Consider two point charges located on the x axis: one charge, q1 14.5 nC is located at z1 1.745 m the second charge, q2 33.5 nC is at the origin (z 0.0000). Part A What is the net force exerted by these two charges on a third charge ga 45.0nC placed between gi and g2 at 1.105 m Your answer may be positive or negative, depending on the direction of the force. Express your answer numerically in newtons to three significant figures. Force on as 3.07.10
Answer #1

The diatance between chrge q1=-14 nC and q2=33.5 nC is 1.745 m .

another charge of 45 nC is placed between them at 1.105 m left from charge q2 .

Distance between 45 nC charge and q2, d=1.745 -1.105 =0.64 m

Force on 45 nC charge due to q1, F1=K (14.5\times10-9\times45\times10-9)/ (0.64)2   (attractive force , direction of force is towards left)

F1=(9\times109)(14.5\times10-9\times45\times10-9)/ (0.64)2

F1=1.43\times10-5 N towards left.

Force on 45 nC charge due to q2,  F2=K (33.5\times10-9\times45\times10-9)/ (1.105)2   (repulsive force ,direction of force is towards left)

F2=(9\times109)(33.5\times10-9\times45\times10-9)/ (1.105)2

F2=1.11\times10-5 N towards left

Net force on 45 nC charge due to both of force is

F=F1+F2   (Both force in same direction)

F=1.43\times10-5 +1.11\times10-5 N

F=2.54\times10-5 N (towards left ) ANS